OpenStudy (anonymous):

Write the expression as either the sine, cosine, or tangent of a single angle. cos(pi/3)cos(pi/5)+sin(pi/3)sin(pi/5)

OpenStudy (anonymous):

I think it probably has to do with the unit circle but I really have no idea what to do for it, so an explanation would be nice also :s

OpenStudy (anonymous):

We know that: \[cos(a\pm b)=cos(a)cos(b)\mp sin(a)sin(b)\] So therefore: \[cos(a)cos(b)+sin(a)sin(b)=cos(a-b)\] So therefore: \[cos\left(\frac{\pi}{3}\right)cos\left(\frac{\pi}{5}\right)+sin\left(\frac{\pi}{3}\right)sin\left(\frac{\pi}{5}\right)=cos\left(\frac{\pi}{3}-\frac{\pi}{5}\right)=cos\left(\frac{2\pi}{15}\right)\]

OpenStudy (anonymous):

How do we know the first part? Is there an identity

OpenStudy (anonymous):

I understand it now, thanks

OpenStudy (anonymous):

Oh lol yeah the first is a identity lol

OpenStudy (anonymous):

http://www.youtube.com/watch?v=zcEMKv5yIYs